OpenStudy (anonymous):

HELP ME PLEASE!! An artifact originally had 16 grams of Carbon-14 present. Determine how many grams of Carbon-14 will be present in 11,430 years.

OpenStudy (anonymous):

i think you are suppose to up A=Pe^rt

OpenStudy (anonymous):

do you know how to solve it @ganeshie8 ??

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

they gave rate of decay ? or they expect us to know half life of C-14 and proceed ?

OpenStudy (anonymous):

well they didnt give rate or say its a half life so i didnt know what to do. in the question before this it asked me to find the half life of it but thats it

ganeshie8 (ganeshie8):

ok. then we can use half-life of C-14 here can u google and tell me halflife of C-14

ganeshie8 (ganeshie8):

wait a sec

ganeshie8 (ganeshie8):

yep.. we need to use halflife here, from that we can compute rate-of-decay

OpenStudy (anonymous):

it said it was 5,730 years

ganeshie8 (ganeshie8):

\(\huge \frac{1}{2} = e^{r*5730}\)

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

great ! take log both sides, and compute "r"

ganeshie8 (ganeshie8):

once you have "r", you can use the formula \(\huge P = Ae^{rt}\)

ganeshie8 (ganeshie8):

can you try ?

OpenStudy (anonymous):

ias a final answer i got 1,591,056

ganeshie8 (ganeshie8):

whats that ? 16 grams decayed to 1.5 million grams ????? hows that possible

ganeshie8 (ganeshie8):

whats the value of "r" ?

OpenStudy (anonymous):

its not decayed its growth

ganeshie8 (ganeshie8):

lol yes hows that possible

OpenStudy (anonymous):

the question is after 11,430 years how many grams of c-14 will there be

ganeshie8 (ganeshie8):

u mean, the question is asking to find 11,430 years back how much quantity of C-14 was present ?

ganeshie8 (ganeshie8):

from the halflife equation, you should get a -ve value for "r"

ganeshie8 (ganeshie8):

if you go back in time, you will put -ve value for time

ganeshie8 (ganeshie8):

then it can become growth. but i guess question is not asking for that

OpenStudy (anonymous):

the question is simply asking how many grams of Carbon - 14 will be present in 11,430 years if the artifact originally had 16 grams

ganeshie8 (ganeshie8):

@agentx5 if you're familiar with this... pls help

OpenStudy (anonymous):

Ok, I see what happened here: First, @ganeshie8 , did you get \(\huge-\frac{ln|2|}{5730}\) for r?

OpenStudy (anonymous):

(I'm assuming you all have the correct half-life written down btw)

OpenStudy (anonymous):

yes 5730 years

ganeshie8 (ganeshie8):

@agentx5 i did get -ve for r. and i know the quantity should decay

ganeshie8 (ganeshie8):

maybe Lulu if you can show ur work how you got 1.59 million it helps

OpenStudy (anonymous):

\(\huge A = 4 \sqrt[191]{2} \approx 4.0145 \ \text{g}\) You must be making a calculation error..

OpenStudy (anonymous):

This makes logical sense too, because it's approximately twice the half-life time so you would expect another half would be gone.

ganeshie8 (ganeshie8):

yeah makes sense to me too :)

OpenStudy (anonymous):

You will have approximately half of or half the Carbon 14 you had 11430 years earlier (16g originally)

OpenStudy (anonymous):

Very rough approximation to know if you're in the right ballpark so to speak: \[16 g \cdot ( \frac{1}{2} * \frac{1}{2}) = 16g \cdot (\frac{1}{4}) = 4\]

OpenStudy (anonymous):

\[\frac{5730 \ yrs}{11430 \ yrs} \approx \frac{1}{2}\]

OpenStudy (anonymous):

If this is a multiple choice exam I would get a little practice in understand the concept and the logic I'm using here to test what the answer should be around before going exact. Unless they professor has two very close numbers on the exam you might even be able to make an educated guess and save yourself a lot of time. You should also know how to work the formulas. The answer with the 191st root above is in fact the perfectly exact answer, and if it was a written exam that's what I would write, with units of grams.

OpenStudy (anonymous):

Make sense @Lulu212 ?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

but this is not an exam its hw

OpenStudy (anonymous):

Example: if my choices were... A) 10g B) 3g C) 4g D) 6g I could just use reasoning to figure out the answer should be damn close to 4 and on the little bit more side.

OpenStudy (anonymous):

Homework is practice for exams ;-)

OpenStudy (anonymous):

And if I was teaching Chemistry 1, you better expect I would have half-life on the exam. It's just one of those fundamental concepts you should know in chemistry.

OpenStudy (anonymous):

but how is it possible to go from 16 to 4 in 11430 if it meant to increase??

OpenStudy (anonymous):

Graph of the amounts present with respect to time. C\(^{14}_6\) decays into Carbon-12. At first quickly, then starting to slow down.